# Anisotropic Field Theory of Wave Transmission Statistics in Disordered Media

### v2

@@ Abstract
Figure 2: Transmission eigenvalue distribution for an absorbing disordered waveguide with $W/\lambda=50.5$, $L/W=1$, $L/\ell_{\rm s}=5$, and $m_{\mathrm{a}}=m_{\mathrm{b}}=1$, shown for different absorption lengths $\ell_{\rm a}$. Numerical distributions (dots) are compared with predictions from the anisotropic field theory (13)–(17) (solid) and its isotropic version based on Eq. (18) (dashed). The inset displays the distribution for $L/\ell_{\rm s}=L/\ell_{\rm a}=0.2$, with the corresponding numerical distribution shown as dotted line.
Interestingly, we find that absorption tends to break the isotropy of $\mathsf{g}$, even in the diffusive regime ($L/\ell_{\rm s}=5$ in Fig. 2), owing to the variation in average path lengths taken by incoming waves based on their angle of incidence. This deviation disappears for $L/\ell_{\rm s}\gg 10$. In the quasiballistic regime, lobes become visible (see inset of Fig. 2), resulting from higher transverse modes which, due to their extended paths in the absorbing medium, contribute to lower transmission. This lobe-splitting effect cannot be captured by an isotropic theory. We also emphasize that no theoretical prediction has been available until now for the experimentally relevant regime of moderate absorption, $\ell_{\rm a}\leq L\ll\xi_{\rm a}$, where $\xi_{\rm a}=\sqrt{\ell_{\rm s}\ell_{\rm a}/2}$ denotes the diffusive absorption length. The approximation proposed in Ref. [46] applies only in the strongly absorbing regime, $L\gg\xi_{\rm a}$, where transmission is minimal.
@@
Finally, we consider an experimental situation where only modes within the input port’s numerical aperture $m_{\mathrm{a}}$ (i.e., $\left\|\boldsymbol{\mathrm{\Omega}}_{\perp}\right\|\leq m_{\mathrm{a}}$) are injected. At the outport port, we assume that all waveguide modes are accessible ($m_{\mathrm{b}}=1$), making the transmission matrix $\mathsf{t}$ rectangular with more rows than columns.
@@
Figure 3: Transmission eigenvalue distribution for a disordered waveguide with $W/\lambda=50.5$, $L/W=1$, $L/\ell_{\rm s}=5$, and $L/\ell_{\rm a}=0$, shown for different input port numerical apertures $m_{\mathrm{a}}$. The output port has maximum numerical aperture ($m_{\mathrm{b}}=1$). Numerical distributions (dots) are compared with the predictions from the field theory (13)–(17) (solid) and FRM theory [49] with $\bar{T}=1/(1+\tfrac{2L}{\pi\ell_{\rm s}})\simeq 0.239$ (dashed).
@@
In conclusion, this Letter introduces an anisotropic field framework for the transmission eigenvalue distribution in disordered media. The core of this approach is the self-consistent matrix transport equation (13), which allows us to predict the correct distribution in the quasiballistic regime—a novel achievement, building on prior work [33, 57, 34] inspired by superconductivity [37, 38, *Larkin1973, *Larkin1975, *Larkin1977, 42]. Our model also rigorously accommodates medium absorption and practical measurement setups. Because of its flexibility, this framework opens up new avenues: it can describe incident beams with finite spatial extent [50, 51], statistics of observables beyond transmission, such as internal field intensity [47, 48], and is adaptable to non-rectilinear geometries [58]. Furthermore, due to its microscopic foundation, it can be extended to capture effects such as anisotropic scattering in biological tissues or the impact of correlated disorder [59]. We anticipate significant breakthroughs in understanding wave propagation in complex media through this approach.
@@ Acknowledgements.
Acknowledgments—The authors thank Romain Pierrat for early-stage discussions and Romain Rescanières for valuable discussions and additional numerical simulations. This research has been supported by the ANR project MARS_light under reference ANR-19-CE30-0026 and by the program “Investissements d’Avenir” launched by the French Government.

### v3

@@ Abstract
Figure 2: Transmission eigenvalue distribution for an absorbing disordered waveguide with $W/\lambda=50.5$, $L/W=1$, $L/\ell=5$, and $m_{\mathrm{a}}=m_{\mathrm{b}}=1$, shown for different absorption lengths $\ell_{\rm a}$. Numerical distributions (dots) [42] are compared with predictions from the RFT (13)–(17) (solid) [41] and its direction averaged version (18) (dashed). The solid gray line corresponds to the RFT prediction without absorption. The inset displays the distribution for $L/\ell=L/\ell_{\rm a}=0.2$, with the corresponding numerical distribution shown as dotted line and the RFT prediction (13)–(17) as solid line.
In Fig. 2, the RFT (13)–(17) (solid lines) is compared to the numerical distributions based on the microscopic wave equation (dots), in a strongly scattering system ($L/\ell=5$). We see that the main effect of absorption is to contract the transmission spectrum toward $T=0$, while preserving the peak near the upper edge of the spectrum, in agreement with the numerical results of Ref. [45]. Interestingly, this contraction effect is very different from the effect of Anderson localization, which removes the peak at $T=1$ without creating a gap [46]. Furthermore, the deviation between the dashed lines in Fig. 2, corresponding to Eq. (18) with $\varepsilon=k/\ell_{\rm a}$, and the solid lines, suggests that absorption tends to break the validity of the diffusion approximation. Such deviations are not observed without absorption (see the $L/\ell=5$ curves in Fig. 1), and they vanish in the deep diffusive regime ($L/\ell\gg 10$). In the quasiballistic regime ($L/\ell=0.2$, inset of Fig. 2), lobes become visible, resulting from higher transverse modes which, due to their extended paths in the absorbing medium, contribute to lower transmission.
@@
Finally, we consider an experimental situation where only modes within the input port’s numerical aperture $m_{\mathrm{a}}$ (i.e., $\left\|\boldsymbol{\mathrm{\Omega}}_{\perp}\right\|\leq m_{\mathrm{a}}$) are injected. At the output port, we assume that all waveguide modes are accessible ($m_{\mathrm{b}}=1$), making the transmission matrix $\mathsf{t}$ rectangular with more rows than columns.
@@
Figure 3: Transmission eigenvalue distribution for a disordered waveguide with $W/\lambda=50.5$, $L/W=1$, $L/\ell=5$, and $L/\ell_{\rm a}=0$, shown for different input port numerical apertures $m_{\mathrm{a}}$. The output port has maximum numerical aperture ($m_{\mathrm{b}}=1$). Numerical distributions (dots) [42] are compared with the predictions from the RFT (13)–(17) (solid) [41] and FRM theory [35] with $\bar{T}=1/(1+\tfrac{2L}{\pi\ell})\simeq 0.239$ (dashed). The solid gray line corresponds to the RFT prediction for $m_{\mathrm{a}}=1$.
@@
In conclusion, we have introduced a field-theoretic framework for the transmission eigenvalue distribution in disordered media. At its core lies a self-consistent matrix transport equation [Eq. (13)], which accurately captures the distribution across both the diffusive and quasiballistic regimes. Beyond this, our model naturally incorporates key experimental conditions, such as absorption and incomplete channel control. Owing to its flexibility, the framework can be extended to describe finite-sized incident beams [36, 37, 47], nonrectilinear geometries [48], and, as we will demonstrate in future work, statistics of observables beyond transmission, such as deposition [49, 50], or internal field intensities resulting from the propagation of shaped waves. Thanks to its microscopic foundation, the theory is also well suited to capturing more complex phenomena, including anisotropic scattering in biological tissues or the effects of correlated disorder [51]. Extending this framework to the regime of Anderson localization is an especially promising direction. We anticipate that this approach will enable significant advances in the theoretical understanding of wave propagation in complex media.
@@ Acknowledgements.
Acknowledgments—The authors thank Romain Pierrat for early-stage discussions and Romain Rescanieres for valuable discussions and additional numerical simulations. This research has been supported by the ANR project MARS_light under reference ANR-19-CE30-0026, a grant from the Simons Foundation (No. 1027116), and the program “Investissements d’Avenir” launched by the French Government.